a women is now 24 years older than her niece in 8 years she will be twice as old as her niece will be then find their present ages.
Found 2 solutions by TimothyLamb, MARIO OWHOLO:
Answer by TimothyLamb(4379)

w = n + 24
w + 8 = 2(n + 8)
---
w + 8 = 2(n + 8)
w + 8 = 2n + 16
n + 24 + 8 = 2n + 16
---
n = 16
w = 40
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by MARIO OWHOLO(4)

Woman=24+niece. Equation 1
Woman+8=2(niece)+8 Equation 2
Woman=x
Niece=y
Y=x-24 Equation 3
X+8=2(x-24)+8
X=2x-48+8-8
3x=48
X=16
Therefore
-y=8

Woman=16yrs old
Niece=8yrs old